The width of a rectangle is 12 cm shorter than the length. The perimeter is 52 cm. What is thw demensions of the rectangle ?
Answer by palanisamy(496)

Let the lenghth = x
Then its width = x-12
Perimeter 2=52
2=52
4x-24=52
4x=52+24
4x=76
x=76/4=19
So, the length = 19
width = 19-12 = 7